OpenStudy (anonymous):

Out of a random sample of 500, 75 like oranges. How much more sample should be taken in order to reach the no 100 who like oranges

OpenStudy (katielong):

500:75 x:100 to get from 75 to 25 you can.. divide by 3 then times by 4 so you do the same with 500 to get x 500/3=166.666... 166.666...x4=668 rounded up that may be it

OpenStudy (anonymous):

hey katie does it work if i do this: 500/75= .15 of ppl like oranges 100-75= 25 more ppl we need to find that like oranges 25/.15= 166?

OpenStudy (katielong):

yehh, nice and thats probably more simple than my answer

OpenStudy (anonymous):

ok perfect thank you so much

OpenStudy (katielong):

no probs :)